OpenStudy (anonymous):

The numbers of beads on 500 handcrafted bead necklaces follow a normal distribution whose mean is 24 beads and standard deviation is 4 beads. Which sentence most closely summarizes the data? a. About 24 necklaces have more than 28 beads. b. About 48 necklaces have more than 28 beads. c. About 80 necklaces have more than 28 beads. d. About 96 necklaces have more than 28 beads.

OpenStudy (anonymous):

First, find out what percent of things will be above the first standard deviation away

OpenStudy (anonymous):

You basically have a \(z\) score of \(1\) here. \[ z=\frac{x-\mu}{\sigma}=\frac{28-24}{4} = 1 \]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=z+score+1 Left tailed will be about \(0.1587\) of all necklaces.

OpenStudy (anonymous):

So bascically \(15.87\%\) of the \(500\) necklaces will be above \(28\) beads. http://www.wolframalpha.com/input/?i=15.87%25+of+500

OpenStudy (anonymous):

I don't understand what that means,

OpenStudy (anonymous):

"that"?